# GRU: Mitigating the Trade-off between Unlearning and Retention for Large Language Models

### v1

@@ 5.2 Retain Data Free
We further consider the retain-data-free settings, where we have no retain data at hand as mentioned in Section 4. As a case study, we test the efficacy of various methods that do not rely on retain data and our TRU. We further include the baseline of task vector (TV) [27] for fair comparison, which is also the key technique that is adopted in our TRU. The experimental results are summarized in Table 1, where we also report metric scores for the model before unlearning (original) as well as for the gold standard model (retrain), which is fine-tuned from scratch without the targeted data. Across baseline methods, it can be observed that the retain-data-free settings is challenging. Only WGA and NPO can demonstrate some ability of reliable unlearning. In contrast, other methods, such as TV and GA, can render the unlearned models completely useless. Furthermore, our TRU exhibits notable improvements over these baselines in both removal and retention, showcasing the broad applications of our unlearning schemes suggested in Eq. (7) even in some more restricted unlearning setups.
5.3 Sensitivity Analyses
Method Metric $\tau$
0.001 0.01 0.1 1.0 2.0 3.0 10 100 N/A
Unlearning Strength. We also report a range of strength stg in Table 4 for our task vector-based unlearning approach. The observed numerical stability in both unlearning and retention performance underscores the robust efficacy of our proposed “rectification" framework.
Method Metric $\gamma$
0.01 0.02 0.05 0.1 0.2 0.3 0.4 0.5 0.6 0.8 0.9 0.99 N/A
Table 4: Hyper-parameter turning of RDF on the TOFU-llama-%5 with different Task Vector strength stg.
stg
Metric 0.50 0.55 0.60 0.65 0.70 0.75 0.80 0.85
FQ $\uparrow$ -9.9550 -7.3407 -6.5526 -4.4360 -2.4323 -1.7263 -0.6065 -0.3389
MU $\uparrow$ 0.5509 0.5251 0.4856 0.4507 0.4211 0.3812 0.3526 0.3233
5.4 Ablation Studies.
Previous works, such as [15], also use gradient clipping (GC) to improve the overall efficacy of unlearning, raising us to ask if our rectification mechanism plays a key role in mitigating the trade-off between removal and retention. In Table 5, we conduct ablation studies on TOFU using Llama-2-7B as the backbone, focusing on the 5% unlearning setup. We compare three scenarios: the original unlearning method without GRU (w/o GRU), the original method enhanced with GC (w/ GC), and the unlearning method that incorporates GRU (w/ GRU). As evident from the results, GRU demonstrates superior scores in terms of both FQ and MU, showing its efficacy in mitigating the trade-off between removal and retention.
Method Component FQ $\uparrow$ MU $\uparrow$
GA w/o GRU -16.9281 0.0000
w/ GC -20.7646 0.0000
w/ GRU -3.2299 0.5149
NPO w/o GRU -10.9105 0.4919
w/ GC -10.9105 0.4970
w/ GRU -10.4275 0.5373
@@ 5.5 More Results in Appendices
Due to space limits, we leave more detailed experimental results and additional analyses in the appendices. For convenience, this section provides a brief overview of these contents: In Appendix B, we offer more comprehensive results for our main experiments on varying benchmarks and metrics, further covering other baselines such as SimNPO [35] and NPO+KL. In Appendix C, we detail the hyper-parameter configurations and outline their recommended setups. Finally, in Appendix D, we include other experimental analyses of GRU.
@@ 6 Conclusion
This paper introduces GRU, a novel and general framework designed to mitigate the inherent trade-off between data removal and retention for LLM unlearning, a critical challenge in this field. Our key insight involves regulating the gradients used for unlearning by projecting them onto the orthogonal complement of directions that negatively affect retention. Thereby, GRU ensures that the unlearning updates minimize their adverse impact on the overall model performance. We offer both theoretical analyses and empirical evidence to demonstrate the effectiveness of our method in mitigating the trade-off between removal and retention, resulting in the overall efficacy of unlearning. However, our method critically relies on the quality of retain data. While TRU can mitigate this issue to some extent, potential biases and distribution shifts therein may still be detrimental. In the future, we will explore ways to pursue reliable LLM unlearning without relying on retain data or their surrogates.
Impact Statements
@@ Appendix A Derivations and Proofs
$&\operatorname*{arg\,min}_{\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}}\hskip 2.0pt&&\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}-\boldsymbol{g}_{\mathrm{u}}^{(t)}\|^{2}\\ &\hskip 1.8pt\text{s.t.}~{}&&\langle\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)},{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\rangle\geq 0\end{aligned},$
@@
$\frac{1}{2}\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}-\boldsymbol{g}_{\mathrm{u}}^{(t)}\|^{2}-\kappa\langle\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)},\,{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\rangle$ (14)
with $\kappa\geq 0$ the Lagrange multiplier. Then, setting the gradients Eq. (14) with respect to $\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}$ to zero, we have
$\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}-\boldsymbol{g}_{\mathrm{u}}^{(t)}-\kappa{\boldsymbol{g}}_{\mathrm{r}}^{(t)}={0}.$ (15)
@@
$\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}=\boldsymbol{g}_{\mathrm{u}}^{(t)}+\kappa{\boldsymbol{g}}_{\mathrm{r}}^{(t)}.$
@@
$\langle\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)},\,{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\rangle=\langle\boldsymbol{g}_{\mathrm{u}}^{(t)}+\kappa\boldsymbol{g}_{\mathrm{r}}^{(t)},\,{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\rangle=\langle\boldsymbol{g}_{\mathrm{u}}^{(t)},\,{\boldsymbol{g}}_{\mathrm{r}}^{(t)})\rangle+\kappa\|{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\|^{2}\geq 0$
@@
$\kappa\geq-\frac{\langle\boldsymbol{g}_{\mathrm{u}}^{(t)},\,{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\rangle}{\|{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\|^{2}}.$
@@
$\kappa=\frac{[-\langle\boldsymbol{g}_{\mathrm{u}}^{(t)},\,{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\rangle]_{+}}{\|{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\|^{2}}=\frac{[\langle\boldsymbol{g}_{\mathrm{u}}^{(t)},\,{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\rangle]_{-}}{\|{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\|^{2}}.$
@@
$\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}=\boldsymbol{g}_{\mathrm{u}}^{(t)}+\frac{[\langle\boldsymbol{g}_{\mathrm{u}}^{(t)},\,{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\rangle]_{-}}{\|{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\|^{2}}\,{\boldsymbol{g}}_{\mathrm{r}}^{(t)}$
@@
We further demonstrate that the GRU causes the magnitudes of the rectified gradients to decrease. To this end, we first decompose the original gradients $\boldsymbol{g}_{\mathrm{u}}^{(t)}$ into two orthogonal components that are parallel and perpendicular to $\boldsymbol{g}_{\mathrm{r}}^{(t)}$, which are
$\boldsymbol{g}_{\mathrm{u}}^{(t)}=\boldsymbol{g}_{\perp}+\boldsymbol{g}_{\parallel}~{}\text{and}~{}\boldsymbol{g}_{\perp}\perp\boldsymbol{g}_{\mathrm{r}}^{(t)}.$
with $\boldsymbol{g}_{\parallel}$ parallel to $\boldsymbol{g}_{\mathrm{r}}^{(t)}$. In this decomposition, $\boldsymbol{g}_{\parallel}$ represents the component of $\boldsymbol{g}_{\mathrm{u}}^{(t)}$ that aligns with $\boldsymbol{g}_{\mathrm{r}}^{(t)}$, whereas $\boldsymbol{g}_{\perp}$ is orthogonal to $\boldsymbol{g}_{\mathrm{r}}^{(t)}$. Then, if $\langle\boldsymbol{g}_{\mathrm{u}}^{(t)},\boldsymbol{g}_{\mathrm{r}}^{(t)}\rangle\geq 0$, no adjustment is needed and we keep $\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}=\boldsymbol{g}_{\mathrm{u}}^{(t)}$. In this case, the norm remains the same as $\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\|=\|\boldsymbol{g}_{\mathrm{u}}^{(t)}\|$. However, when $\langle\boldsymbol{g}_{\mathrm{u}}^{(t)},\boldsymbol{g}_{\mathrm{r}}^{(t)}\rangle<0$, $\boldsymbol{g}_{\parallel}$ represents a negatively aligned component with respect to $\boldsymbol{g}_{\mathrm{r}}^{(t)}$. The correction term in Eq. (8) removes this negative parallel portion, thereby setting $\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}=\boldsymbol{g}_{\perp}$. Since $\boldsymbol{g}_{\mathrm{u}}^{(t)}=\boldsymbol{g}_{\perp}+\boldsymbol{g}_{\parallel}$, we have $\|\boldsymbol{g}_{\mathrm{u}}^{(t)}\|^{2}=\|\boldsymbol{g}_{\perp}\|^{2}+\|\boldsymbol{g}_{\parallel}\|^{2}$. When the parallel component is negative relative to $\boldsymbol{g}_{\mathrm{r}}^{(t)}$, its removal decreases the overall norm. Thus, it is easy to conclude that, after rectification, we have
$\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\|=\|\boldsymbol{g}_{\perp}\|<\|\boldsymbol{g}_{\mathrm{u}}^{(t)}\|.$
This difference in magnitudes is influenced by the angles between $\boldsymbol{g}_{\mathrm{u}}^{(t)}$ and $\boldsymbol{g}_{\mathrm{r}}^{(t)}$. Overall, as these angles widen beyond $90^{\circ}$, the magnitudes of the negative component $\boldsymbol{g}_{\parallel}$ increases. It indicates that a greater portion of this components should be removed to fulfill the constraint specified in Eq. (7). In an extreme case where the angles approach $180^{\circ}$, nearly the entire $\boldsymbol{g}_{\mathrm{u}}^{(t)}$ is inverted relative to $\boldsymbol{g}_{\mathrm{r}}^{(t)}$. It implies a substantial reduction in magnitudes of $\boldsymbol{g}_{\mathrm{u}}^{(t)}$ to approach 0 after adjustment.
@@
To simply our symbology, we use $\mathcal{L}(\boldsymbol{\theta})$ and $\mathcal{R}(\boldsymbol{\theta})$ to replace $\mathcal{L}(\mathcal{D}_{\mathrm{u}};\boldsymbol{\theta})$ and $\mathcal{R}(\mathcal{D}_{\mathrm{r}};\boldsymbol{\theta})$ if raising no confusion.
@@
$\mathcal{L}(\boldsymbol{\theta}^{(t)}-\texttt{lr}\,\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)})$ $\;\leq\;\mathcal{L}(\boldsymbol{\theta}^{(t)})\;-\;\texttt{lr}\,\boldsymbol{g}_{\mathrm{u}}^{(t)\top}\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\;+\;\frac{L\,\texttt{lr}^{2}}{2}\,\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\|^{2}.$
@@
$\Delta^{(t)}\;:=\;-\,\texttt{lr}\;\boldsymbol{g}_{\mathrm{u}}^{(t)\top}\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\;+\;\frac{L\,\texttt{lr}^{2}}{2}\,\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\|^{2}.$
Then, if $\Delta^{(t)}<0$, we have a strict decrease, i.e., $\mathcal{L}(\boldsymbol{\theta}^{(t+1)})<\mathcal{L}(\boldsymbol{\theta}^{(t)})$, and thus we can complete our proof. As observed, $\Delta^{(t)}$ consists of two terms:
  • Linear term: Since $\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}$ is a projection of $\boldsymbol{g}_{\mathrm{u}}^{(t)}$ that does not invert the direction, we have $\boldsymbol{g}_{\mathrm{u}}^{(t)\top}\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\geq 0$.
  • Quadratic term: Since the norm of $\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}$ is bounded by $\|\boldsymbol{g}_{\mathrm{u}}^{(t)}\|$, we have $\frac{L\texttt{lr}^{2}}{2}\,\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\|^{2}\;\;\leq\;\;\frac{L\texttt{lr}^{2}}{2}\,\|\boldsymbol{g}_{\mathrm{u}}^{(t)}\|^{2}$.
@@
$\mathcal{L}(\boldsymbol{\theta}^{(t+1)})\;<\;\mathcal{L}(\boldsymbol{\theta}^{(t)}).$
Thus, we obtain a strict descent unless we encounter a degenerate scenario. Specifically, if $\boldsymbol{g}_{\mathrm{u}}^{(t)}$ happens to be exactly reversed with respect to the retain gradients $\boldsymbol{g}_{\mathrm{r}}^{(t)}$, i.e. their angle is $180^{\circ}$ and $\cos(\boldsymbol{g}_{\mathrm{u}}^{(t)},\boldsymbol{g}_{\mathrm{r}}^{(t)})=-1$, then, after rectification, one obtains $\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}={0}$. In this case, we have $\boldsymbol{\theta}^{(t+1)}\;=\;\boldsymbol{\theta}^{(t)}$, which makes no further decrease. Thus, we complete the proof. ∎
@@
For any smooth and differentiable loss $\mathcal{R}$, its $q$-Curvature with respect to some gradients $\boldsymbol{g}$ is defined as
@@
Recall that at the $t$-th iteration, the original updating rule without GRU is $\boldsymbol{\theta}_{\mathrm{u}}^{(t+1)}=\boldsymbol{\theta}_{\mathrm{u}}^{(t)}-\texttt{lr}\,\boldsymbol{g}_{\mathrm{u}}^{(t)}$. Additionally, according to the integral form of Taylor theorem, for any $\alpha\in[0,1]$, we have
@@
Since we assume $\mathfrak{H}_{\texttt{lr}}(\mathcal{R};\,\boldsymbol{g}_{\mathrm{u}}^{(t)})~{}\geq~{}\ell\,\|\boldsymbol{g}_{\mathrm{u}}^{(t)}\|^{2}$, we have
@@
$\mathcal{R}\!(\boldsymbol{\theta}^{(t)}-\texttt{lr}\,\boldsymbol{g}_{\mathrm{u}}^{(t)})~{}\geq~{}\mathcal{R}\!(\boldsymbol{\theta}^{(t)})\;-\;\texttt{lr}\,\langle\boldsymbol{g}_{\mathrm{r}}^{(t)},\,\boldsymbol{g}_{\mathrm{u}}^{(t)}\rangle\;+\;\frac{\ell\,\texttt{lr}^{2}}{2}\,\|\boldsymbol{g}_{\mathrm{u}}^{(t)}\|^{2},$ (17)
which establishes the lower bound for $\mathcal{R}\!(\boldsymbol{\theta}_{\mathrm{u}}^{(t+1)})=\mathcal{R}\!(\boldsymbol{\theta}^{(t)}-\texttt{lr}\,\boldsymbol{g}_{\mathrm{u}}^{(t)}).$ For the rectified updating rule with GRU, due to the $L$-smoothness, we have
$\mathcal{R}\!(\boldsymbol{\theta}_{\mathrm{gru}}^{(t+1)})\;\;\leq\;\;\mathcal{R}\!(\boldsymbol{\theta}^{(t)})\;-\;\texttt{lr}\,\langle\boldsymbol{g}_{\mathrm{r}}^{(t)},\;\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\rangle\;+\;\frac{L\,\texttt{lr}^{2}}{2}\,\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\|^{2}.$ (18)

### v3

@@ 5.2 Retain Data Free
We further consider the retain-data-free settings, where we have no retain data at hand as mentioned in Section 4. As a case study, we test the efficacy of various methods that do not rely on retain data and our TRU. We further include the baseline of task vector (TV) (Ilharco et al., 2022) for fair comparison, which is also the key technique that is adopted in our TRU. The experimental results are summarized in Table 1, where we also report metric scores for the model before unlearning (original) as well as for the gold standard model (retrain), which is fine-tuned from scratch without the targeted data. Across baseline methods, it can be observed that the retain-data-free settings is challenging. Only WGA and NPO can demonstrate some ability of reliable unlearning. In contrast, other methods, such as TV and GA, can render the unlearned models completely useless. Furthermore, our TRU exhibits notable improvements over these baselines in both removal and retention, showcasing the broad applications of our unlearning schemes suggested in Eq. (7) even in some more restricted unlearning setups.
@@ 5.3 More Results in Appendices
Due to space limit, we leave more detailed experimental results and additional analyses to the appendices. For convenience, this section provides a brief overview of these contents: In Appendix B, we offer more comprehensive results for our main experiments on varying benchmarks and metrics, further covering other baselines such as SimNPO (Fan et al., 2024) and NPO+KL. Additionally, we include a comparison with RMU, the method proposed alongside WMDP, specifically evaluated on the WMDP benchmark together with its combination with GRU (see Appendix C for results and discussion). In Appendix D, we perform a hyper-parameter sensitivity analysis and outline their recommended setups. Finally, in Appendix E, we include our ablation studies and other experimental analyses. Finally, we provide a more practically Meaningful and fairer comparative analysis in Section F.
@@ 6 Conclusion
This paper introduces GRU, a novel and general framework designed to mitigate the inherent trade-off between data removal and retention for LLM unlearning, a critical challenge in this field. Our key insight involves regulating the gradients used for unlearning by projecting them onto the orthogonal complement of directions that negatively affect retention. Thereby, GRU ensures that the unlearning updates minimize their adverse impact on the overall model performance. We offer both theoretical analyses and empirical evidence to demonstrate the effectiveness of our method in mitigating the trade-off between removal and retention, resulting in overall efficacy of unlearning. However, our method critically relies on the quality of retain data. While TRU can mitigate this issue to some extent, potential biases and distribution shifts therein may still be detrimental. In the future, we will explore ways to pursue reliable LLM unlearning without relying on retain data or their surrogates.
Impact Statement
@@
Acknowledgements
YW, QZW, and BH were supported by HKBU Faculty Niche Research Areas No. RC-FNRA-IG/22-23/SCI/04 and HKBU CSD Departmental Incentive Scheme. FL was supported by the Australian Research Council (ARC) with grant number DE240101089, LP240100101, DP230101540 and the NSF&CSIRO Responsible AI program with grant number 2303037. The authors also would like to express their sincere gratitude to the anonymous reviewers and the area chairs for their thorough review and constructive feedback. Their insightful comments and valuable suggestions have significantly enhanced the quality and clarity of this manuscript. We deeply appreciate their effort in helping us improve our work.
@@ Appendix A Derivations and Proofs
$&\operatorname*{arg\,min}_{\tilde{\bm{g}}_{\mathrm{u}}^{(t)}}\hskip 2.0pt&&\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}-\bm{g}_{\mathrm{u}}^{(t)}\|^{2}\\ &\hskip 1.8pt\text{s.t.}~{}&&\langle\tilde{\bm{g}}_{\mathrm{u}}^{(t)},{\bm{g}}_{\mathrm{r}}^{(t)}\rangle\geq 0\end{aligned},$
@@
$\frac{1}{2}\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}-\bm{g}_{\mathrm{u}}^{(t)}\|^{2}-\kappa\langle\tilde{\bm{g}}_{\mathrm{u}}^{(t)},\,{\bm{g}}_{\mathrm{r}}^{(t)}\rangle$ (14)
with $\kappa\geq 0$ the Lagrange multiplier. Then, setting the gradients Eq. (14) with respect to $\tilde{\bm{g}}_{\mathrm{u}}^{(t)}$ to zero, we have
$\tilde{\bm{g}}_{\mathrm{u}}^{(t)}-\bm{g}_{\mathrm{u}}^{(t)}-\kappa{\bm{g}}_{\mathrm{r}}^{(t)}={0}.$ (15)
@@
$\tilde{\bm{g}}_{\mathrm{u}}^{(t)}=\bm{g}_{\mathrm{u}}^{(t)}+\kappa{\bm{g}}_{\mathrm{r}}^{(t)}.$
@@
$\langle\tilde{\bm{g}}_{\mathrm{u}}^{(t)},\,{\bm{g}}_{\mathrm{r}}^{(t)}\rangle=\langle\bm{g}_{\mathrm{u}}^{(t)}+\kappa\bm{g}_{\mathrm{r}}^{(t)},\,{\bm{g}}_{\mathrm{r}}^{(t)}\rangle=\langle\bm{g}_{\mathrm{u}}^{(t)},\,{\bm{g}}_{\mathrm{r}}^{(t)})\rangle+\kappa\|{\bm{g}}_{\mathrm{r}}^{(t)}\|^{2}\geq 0$
@@
$\kappa\geq-\frac{\langle\bm{g}_{\mathrm{u}}^{(t)},\,{\bm{g}}_{\mathrm{r}}^{(t)}\rangle}{\|{\bm{g}}_{\mathrm{r}}^{(t)}\|^{2}}.$
@@
$\kappa=\frac{[-\langle\bm{g}_{\mathrm{u}}^{(t)},\,{\bm{g}}_{\mathrm{r}}^{(t)}\rangle]_{+}}{\|{\bm{g}}_{\mathrm{r}}^{(t)}\|^{2}}=\frac{[\langle\bm{g}_{\mathrm{u}}^{(t)},\,{\bm{g}}_{\mathrm{r}}^{(t)}\rangle]_{-}}{\|{\bm{g}}_{\mathrm{r}}^{(t)}\|^{2}}.$
@@
$\tilde{\bm{g}}_{\mathrm{u}}^{(t)}=\bm{g}_{\mathrm{u}}^{(t)}+\frac{[\langle\bm{g}_{\mathrm{u}}^{(t)},\,{\bm{g}}_{\mathrm{r}}^{(t)}\rangle]_{-}}{\|{\bm{g}}_{\mathrm{r}}^{(t)}\|^{2}}\,{\bm{g}}_{\mathrm{r}}^{(t)}$
@@
We further demonstrate that the GRU causes the magnitudes of the rectified gradients to decrease. To this end, we first decompose the original gradients $\bm{g}_{\mathrm{u}}^{(t)}$ into two orthogonal components that are parallel and perpendicular to $\bm{g}_{\mathrm{r}}^{(t)}$, which are
$\bm{g}_{\mathrm{u}}^{(t)}=\bm{g}_{\perp}+\bm{g}_{\parallel}~{}\text{and}~{}\bm{g}_{\perp}\perp\bm{g}_{\mathrm{r}}^{(t)}.$
with $\bm{g}_{\parallel}$ parallel to $\bm{g}_{\mathrm{r}}^{(t)}$. In this decomposition, $\bm{g}_{\parallel}$ represents the component of $\bm{g}_{\mathrm{u}}^{(t)}$ that aligns with $\bm{g}_{\mathrm{r}}^{(t)}$, whereas $\bm{g}_{\perp}$ is orthogonal to $\bm{g}_{\mathrm{r}}^{(t)}$. Then, if $\langle\bm{g}_{\mathrm{u}}^{(t)},\bm{g}_{\mathrm{r}}^{(t)}\rangle\geq 0$, no adjustment is needed and we keep $\tilde{\bm{g}}_{\mathrm{u}}^{(t)}=\bm{g}_{\mathrm{u}}^{(t)}$. In this case, the norm remains the same as $\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\|=\|\bm{g}_{\mathrm{u}}^{(t)}\|$. However, when $\langle\bm{g}_{\mathrm{u}}^{(t)},\bm{g}_{\mathrm{r}}^{(t)}\rangle<0$, $\bm{g}_{\parallel}$ represents a negatively aligned component with respect to $\bm{g}_{\mathrm{r}}^{(t)}$. The correction term in Eq. (8) removes this negative parallel portion, thereby setting $\tilde{\bm{g}}_{\mathrm{u}}^{(t)}=\bm{g}_{\perp}$. Since $\bm{g}_{\mathrm{u}}^{(t)}=\bm{g}_{\perp}+\bm{g}_{\parallel}$, we have $\|\bm{g}_{\mathrm{u}}^{(t)}\|^{2}=\|\bm{g}_{\perp}\|^{2}+\|\bm{g}_{\parallel}\|^{2}$. When the parallel component is negative relative to $\bm{g}_{\mathrm{r}}^{(t)}$, its removal decreases the overall norm. Thus, it is easy to conclude that, after rectification, we have
$\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\|=\|\bm{g}_{\perp}\|<\|\bm{g}_{\mathrm{u}}^{(t)}\|.$
This difference in magnitudes is influenced by the angles between $\bm{g}_{\mathrm{u}}^{(t)}$ and $\bm{g}_{\mathrm{r}}^{(t)}$. Overall, as these angles widen beyond $90^{\circ}$, the magnitudes of the negative component $\bm{g}_{\parallel}$ increases. It indicates that a greater portion of this components should be removed to fulfill the constraint specified in Eq. (7). In an extreme case where the angles approach $180^{\circ}$, nearly the entire $\bm{g}_{\mathrm{u}}^{(t)}$ is inverted relative to $\bm{g}_{\mathrm{r}}^{(t)}$. It implies a substantial reduction in magnitudes of $\bm{g}_{\mathrm{u}}^{(t)}$ to approach 0 after adjustment.
@@
To simply our symbology, we use $\mathcal{L}(\bm{\theta})$ and $\mathcal{R}(\bm{\theta})$ to replace $\mathcal{L}(\mathcal{D}_{\mathrm{u}};\bm{\theta})$ and $\mathcal{R}(\mathcal{D}_{\mathrm{r}};\bm{\theta})$ if raising no confusion.
@@
$\mathcal{L}(\bm{\theta}^{(t)}-\texttt{lr}\,\tilde{\bm{g}}_{\mathrm{u}}^{(t)})$ $\;\leq\;\mathcal{L}(\bm{\theta}^{(t)})\;-\;\texttt{lr}\,\bm{g}_{\mathrm{u}}^{(t)\top}\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\;+\;\frac{L\,\texttt{lr}^{2}}{2}\,\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\|^{2}.$
@@
$\Delta^{(t)}\;:=\;-\,\texttt{lr}\;\bm{g}_{\mathrm{u}}^{(t)\top}\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\;+\;\frac{L\,\texttt{lr}^{2}}{2}\,\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\|^{2}.$
Then, if $\Delta^{(t)}<0$, we have a strict decrease, i.e., $\mathcal{L}(\bm{\theta}^{(t+1)})<\mathcal{L}(\bm{\theta}^{(t)})$, and thus we can complete our proof. As observed, $\Delta^{(t)}$ consists of two terms:
  • Linear term: Since $\tilde{\bm{g}}_{\mathrm{u}}^{(t)}$ is a projection of $\bm{g}_{\mathrm{u}}^{(t)}$ that does not invert the direction, we have $\bm{g}_{\mathrm{u}}^{(t)\top}\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\geq 0$.
  • Quadratic term: Since the norm of $\tilde{\bm{g}}_{\mathrm{u}}^{(t)}$ is bounded by $\|\bm{g}_{\mathrm{u}}^{(t)}\|$, we have $\frac{L\texttt{lr}^{2}}{2}\,\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\|^{2}\;\;\leq\;\;\frac{L\texttt{lr}^{2}}{2}\,\|\bm{g}_{\mathrm{u}}^{(t)}\|^{2}$.
@@
$\mathcal{L}(\bm{\theta}^{(t+1)})\;<\;\mathcal{L}(\bm{\theta}^{(t)}).$
Thus, we obtain a strict descent unless we encounter a degenerate scenario. Specifically, if $\bm{g}_{\mathrm{u}}^{(t)}$ happens to be exactly reversed with respect to the retain gradients $\bm{g}_{\mathrm{r}}^{(t)}$, i.e. their angle is $180^{\circ}$ and $\cos(\bm{g}_{\mathrm{u}}^{(t)},\bm{g}_{\mathrm{r}}^{(t)})=-1$, then, after rectification, one obtains $\tilde{\bm{g}}_{\mathrm{u}}^{(t)}={0}$. In this case, we have $\bm{\theta}^{(t+1)}\;=\;\bm{\theta}^{(t)}$, which makes no further decrease. Thus, we complete the proof. ∎
@@
For any smooth and differentiable loss $\mathcal{R}$, its $q$-Curvature with respect to some gradients $\bm{g}$ is defined as
@@
Recall that at the $t$-th iteration, the original updating rule without GRU is $\bm{\theta}_{\mathrm{u}}^{(t+1)}=\bm{\theta}_{\mathrm{u}}^{(t)}-\texttt{lr}\,\bm{g}_{\mathrm{u}}^{(t)}$. Additionally, according to the integral form of Taylor theorem, for any $\alpha\in[0,1]$, we have
@@
Since we assume $\mathfrak{H}_{\texttt{lr}}(\mathcal{R};\,\bm{g}_{\mathrm{u}}^{(t)})~{}\geq~{}\ell\,\|\bm{g}_{\mathrm{u}}^{(t)}\|^{2}$, we have
@@
$\mathcal{R}\!(\bm{\theta}^{(t)}-\texttt{lr}\,\bm{g}_{\mathrm{u}}^{(t)})~{}\geq~{}\mathcal{R}\!(\bm{\theta}^{(t)})\;-\;\texttt{lr}\,\langle\bm{g}_{\mathrm{r}}^{(t)},\,\bm{g}_{\mathrm{u}}^{(t)}\rangle\;+\;\frac{\ell\,\texttt{lr}^{2}}{2}\,\|\bm{g}_{\mathrm{u}}^{(t)}\|^{2},$ (17)
which establishes the lower bound for $\mathcal{R}\!(\bm{\theta}_{\mathrm{u}}^{(t+1)})=\mathcal{R}\!(\bm{\theta}^{(t)}-\texttt{lr}\,\bm{g}_{\mathrm{u}}^{(t)}).$ For the rectified updating rule with GRU, due to the $L$-smoothness, we have
$\mathcal{R}\!(\bm{\theta}_{\mathrm{gru}}^{(t+1)})\;\;\leq\;\;\mathcal{R}\!(\bm{\theta}^{(t)})\;-\;\texttt{lr}\,\langle\bm{g}_{\mathrm{r}}^{(t)},\;\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\rangle\;+\;\frac{L\,\texttt{lr}^{2}}{2}\,\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\|^{2}.$ (18)